# Automatic gradient descent with generalized Newton’s method

### v1

@@ C.4 Initial learning rate
GeN optimizer and existing parameter-free methods such as D-adaptation [13] and DoG [29] still need someone hyperparameters. To be specific, D-adaptation needs an initial D estimate and a growth rate $\approx 1$ if the training is unstable; DoG needs to carefully select $r_{\epsilon}$ to compute the initial distance: too large values could make DoG diverge, while too small values cannot optimize the model. In our case, GeN only needs one hyperparameter – the initial learning rate $\eta_{0}$, which can be determined with nearly zero effort of manual tuning. We now introduce two methods to set $\eta_{0}$.
@@ D.1 Heuristic learning rate scheduler
Heuristic Constant $c_{0}$ $c_{0}$ 1 [56]
Cosine decay $c_{0}(1+\cos{(t\pi/T)})/2$ $c_{0}$ 1 [40, 53]
@@
WarmpUp $c_{\min}+\left(\frac{t(c_{0}-c_{\min})}{pT}\right)$ $c_{0},c_{\min},p$ 3 [21]
@@
In deep learning, state-of-the-art performance is usually obtained by a multi-hyparameter scheduler: for instance, LLAMA [65] combines linear WarmUp and cosine decay (not decaying to 0 but to a minimum value); ResNet [24] uses the stepwise scheduler (starting at $\eta=0.1$ and decay by 10 times at 32k and 48k iterations) with 5 hyperparameters. While tuning multiple hyperparameters may be feasible for some models, it is generally expensive for large models (say over 1B parameters).
@@ D.2 Automatic learning rate
A number of automatic or parameter-free learning rate methods have attempted to dynamically adjust the learning rate throughout the training. D-adaptation [13] and Prodigy [46] both improve on AdaGrad by characterizing $||\bm{w}_{0}-\bm{w}_{*}||$, where $\bm{w}_{*}$ is the global minimum. However, this motivation is not well-defined in deep learning as the minimum is not unique. In fact, for all our computer vision experiments, D-adapt SGD with even a small magnitude of weight decay diverges. Also Prodigy only implements to Adam, hence not future-proof for more advanced algorithms such as Sophia.
@@
DoG [29] and DoWG [32] are parameter-free dynamic SGD scheduler with convergence guarantees for stochastic convex optimization. Additionally, these methods only work for SGD but not for the adaptive optimizers like Adam, and the algorithms requires to store the past iterates (because the output is the weighted $\sum_{t}c_{t}\bm{w}_{t}$ in Equation 2 of [29] and Theorem 1 of [32], instead of the last iterate $\bm{w}_{T}$). As the authors admit in their Github: ‘DoG must be combined with iterate averaging’, which adds to the memory cost that may be unacceptable for large models. In practice, we have observed these methods to be very unstable and inaccurate, e.g. on toy tasks like fine-tuning CIFAR10/100.

### v3

@@ C.4 Initial learning rate
GeN optimizer and existing parameter-free methods such as D-adaptation Defazio & Mishchenko (2023) and DoG Ivgi et al. (2023) still need someone hyperparameters. To be specific, D-adaptation needs an initial D estimate and a growth rate $\approx 1$ if the training is unstable; DoG needs to carefully select $r_{\epsilon}$ to compute the initial distance: too large values could make DoG diverge, while too small values cannot optimize the model. In our case, GeN only needs one hyperparameter – the initial learning rate $\eta_{0}$, which can be determined with nearly zero effort of manual tuning. We now introduce two methods to set $\eta_{0}$.
@@ C.4.2 Automatic search
C.5 Exact algorithm for GeN
We give an algorithm that computes (3.2) precisely using higher order differentiation. In contrast to Algorithm 1 that approximately computes GeN, this algorithm uses the second-order back-progation, which is generally inefficient and incompatible with large-scale distributed systems due to lack of support.
Algorithm 2 Generalized Newton’s optimizers with higher order differentiation (GeN-HOD)
1:for $t\in 1,\cdots,T$ do
2:     Compute loss $L(\bm{w}_{t})$ by the standard forward pass
3:     Compute vanilla gradient $\mathbf{g}_{t}$ by the first-order back-propagation from $L(\bm{w}_{t})\in\mathbb{R}$
4:     Compute modified gradient $\mathbf{g}_{t}^{\text{optim}}(\bm{w}_{t})$
5:     Compute the Hessian-vector product $\mathbf{H}_{t}\mathbf{g}_{t}^{\text{optim}}$ by the second-order back-propagation from $\mathbf{g}_{t}^{\top}\mathbf{g}_{t}^{\text{optim}}\in\mathbb{R}$
6:     Derive $\mathbf{G}_{t}^{\top}\mathbf{g}_{t}^{\text{optim}}$ and $(\mathbf{g}_{t}^{\text{optim}})^{\top}\mathbf{H}_{t}\mathbf{g}_{t}^{\text{optim}}$ by multiplication
7:     Derive the optimal learning rate $\eta_{t}=\frac{\mathbf{G}_{t}^{\top}\mathbf{g}_{t}^{\text{optim}}}{(\mathbf{g}_{t}^{\text{optim}})^{\top}\mathbf{H}_{t}\mathbf{g}_{t}^{\text{optim}}}$ by (3.1)
8:     Update $\bm{w}_{t+1}=\bm{w}_{t}-\eta_{t}\mathbf{g}_{t}$
C.6 Preliminary results on pre-training
Following https://github.com/kuangliu/pytorch-cifar.git, we experimented on CIFAR10 training from scratch on ResNet18. We replace SGD with AdamW at learning rate 1e-3 but the rest settings are the same as in the repository. We use $\Phi=4$ for GeN. For D-adaptation, we use their AdamW version for a fair comparison.
Figure 15: CIFAR10 pre-training on ResNet18 for 15 epochs.
C.7 Dependence on total steps
Total steps $T$ is an important factor in existing learning rate schedules (e.g. linear and consine decay in Table 8): larger $T$ allows longer large-learning-rate-training, while smaller $T$ requires early learning rate decay. Our GeN in (3.1) is originally proposed as $T$-independent. Yet it may be slightly improved with a dependence of $T$, especially if $T$ is relatively small:
$\eta_{\text{GeN}}^{*}(\mathbf{g}_{t}^{\text{optim}})\longleftarrow\eta_{\text{GeN}}^{*}(\mathbf{g}_{t}^{\text{optim}})\cdot(1-t/T)$ (C.1)
where the last term is a hyperparameter-free factor, decaying from 1 to 0. The form of this factor is flexible, e.g. instead of linear decay, we can also use cosine decay like in D-adaptation Defazio & Mishchenko (2023) and Prodigy Mishchenko & Defazio , which denote such factor as $\gamma_{k}$.
We have used this variant of GeN in the experiments of Places365, INat2021 and NLU.
@@ D.1 Heuristic learning rate scheduler
Heuristic Constant $c_{0}$ $c_{0}$ 1 Raffel et al. (2020)
Cosine decay $c_{0}(1+\cos{(t\pi/T)})/2$ $c_{0}$ 1 Loshchilov & Hutter (2016); Radford et al. (2021)
@@
WarmpUp $c_{\min}+\left(\frac{t(c_{0}-c_{\min})}{pT}\right)$ $c_{0},c_{\min},p$ 3 Goyal et al. (2017)
@@
In deep learning, state-of-the-art performance is usually obtained by a multi-hyparameter scheduler: for instance, LLAMA Touvron et al. (2023a) combines linear WarmUp and cosine decay (not decaying to 0 but to a minimum value); ResNet He et al. (2016) uses the stepwise scheduler (starting at $\eta=0.1$ and decay by 10 times at 32k and 48k iterations) with 5 hyperparameters. While tuning multiple hyperparameters may be feasible for some models, it is generally expensive for large models (say over 1B parameters).
@@ D.2 Automatic learning rate
A number of automatic or parameter-free learning rate methods have attempted to dynamically adjust the learning rate throughout the training. D-adaptation Defazio & Mishchenko (2023) and Prodigy Mishchenko & Defazio both improve on AdaGrad by characterizing $||\bm{w}_{0}-\bm{w}_{*}||$, where $\bm{w}_{*}$ is the global minimum. However, this motivation is not well-defined in deep learning as the minimum is not unique. In fact, for all our computer vision experiments, D-adapt SGD with even a small magnitude of weight decay diverges. Also Prodigy only implements to Adam, hence not future-proof for more advanced algorithms such as Sophia.
@@
DoG Ivgi et al. (2023) and DoWG Khaled et al. (2023) are parameter-free dynamic SGD scheduler with convergence guarantees for stochastic convex optimization. Additionally, these methods only work for SGD but not for the adaptive optimizers like Adam, and the algorithms requires to store the past iterates (because the output is the weighted $\sum_{t}c_{t}\bm{w}_{t}$ in Equation 2 of Ivgi et al. (2023) and Theorem 1 of Khaled et al. (2023), instead of the last iterate $\bm{w}_{T}$). As the authors admit in their Github: ‘DoG must be combined with iterate averaging’, which adds to the memory cost that may be unacceptable for large models. In practice, we have observed these methods to be very unstable and inaccurate, e.g. on toy tasks like fine-tuning CIFAR10/100.
D.3 Second-order methods
There is a long list of second-order or quasi-second-order methods in the optimization literature, including but not limited to stochastic Newton methods Byrd et al. (2016); Wang et al. (2017); Lucchi et al. (2015); Schraudolph et al. (2007), sketching methods Yuan et al. (2022); Luo et al. (2016) and diagonal Hessian methods in Section 3.2. We notice these methods are difficult to implement in distributed learning at large scale, because the Hessian information (or Hessian-vector product) is generally not supported by auto-differentiation or too expensive to store for large models.
D.4 Prior works similar to GeN
We elaborate multiple prior works that also present GeN-SGD in (3.1) and leverage the second-order Taylor approximation in (2.3). However, these methods are limited to $\mathbf{g}^{\text{optim}}=\mathbf{g}^{\text{SGD}}$ and not generalize to pre-conditioned gradients like in AdamW. As a consequence, the role of pre-conditioning of gradient is under-investigated, and the connection to the generalized inverse in (3.3) and to Newton’s method in Proposition 3.3 is lacking.
Importantly, a missing piece in these works is to validate the second-order Taylor approximation of $L(\bm{w}-\eta\mathbf{g})$. Our empirical visualization in Figure 2 serves as a necessary validation, without which the method and algorithms are not justified in the first place [^7] In fact, Figure 1 in Fu & Wu (2024) contradicts their own motivation in Equation (4), which is the Taylor expansion at $\eta=0$, since there is no quadratic curve between $\eta\in[0,\alpha]$..
Experiment-wise, prior works are focused on computer vision and convolutional neural networks, whereas we have extensively tested various model architectures (transformers, GAN, recommendation systems) over many tasks like language modeling.
LQA
The closest work to ours is LQA Zhu et al. (2021) which proposes (3.5).
At the method level, this method (termed LQA-SGD) is equivalent to GeN-SGD but it is not extended to general optimizers: in their Figure 1-4, where the red curves in each figure are exactly the same, LQA-SGD is compared to SGD, SGD-momentum, and Adam; but there is no LQA-Adam. Additionally, the convergence speed of LQA-SGD (or GeN-SGD) is missing, whereas we provide an analysis under the convex Lipschitz regime in Appendix E.
At the algorithm level, LQA relies on estimating the losses at $\{L(\bm{w}-\eta\mathbf{g}),L(\bm{w}),L(\bm{w}+\eta\mathbf{g})\}$:
$\eta_{\text{LQA}}^{*}=\frac{\eta}{2}\frac{L(\bm{w}+\eta\mathbf{g})-L(\bm{w}-\eta\mathbf{g})}{L(\bm{w}+\eta\mathbf{g})-2L(\bm{w})+L(\bm{w}-\eta\mathbf{g})}$
This fixed choice of learning rates gives a closed form but could limit the general applicability when one considers other learning rates (see our discussion below (3.5)). The estimation error of LQA was not analyzed. In contrast, our Proposition 3.6 indicates the benefit of a large batch size which is empirically verified in Figure 3. Additionally, we adopt critical techniques in Remark 3.5 like smoothing and lazy frequency, in order to stabilize the training and enhance the computational efficiency, which are not presented in LQA.
QLABGrad
While LQA and GeN requires 2 extra forward passes, QLABGrad Fu & Wu (2024) only requires 1 extra forward passes by restricting to SGD (hence incompatible to AdamW and other adaptive optimizers).
$\eta^{*}_{\text{QLABGrad}}=\frac{\eta^{2}}{2}\frac{\|\nabla L\|^{2}}{L(\bm{w}-\eta\mathbf{g})-L(\bm{w})+\eta\|\nabla L\|^{2}}.$
QLABGrad provides a convergence analysis in terms of the gradient norm, i.e. $||\nabla L_{t}||\to 0$, whereas our convergence analysis in Appendix E is in terms of $L_{t}\to 0$.
ADLER
ADLER Balboni & Bacciu (2023) uses Hessian-vector product to compute $\mathbf{H}_{t}\mathbf{g}_{t}$ (viewed as a sub-case of Algorithm 2). We note that Hessian-vector product is not only slow, but also not supported yet in distributed learning systems. Nevertheless, ADLER indeed implements (3.2) without approximating $\eta^{*}_{\text{GeN}}$, whereas the estimation error must be considered for other algorithms since $\eta^{*}_{\text{LQA}}\approx\eta^{*}_{\text{GeN}}\approx\eta^{*}_{\text{QLABGrad}}$.
D.5 Applying GeN to more problems
We expect general applicability of GeN to problems beyond the formulation of (2.1). Examples include differentially private optimization Liu & Bu , multi-task optimization Jin et al. (2025), gradient ascent, projected/proximal gradient descent, reinforcement learning, etc.
Appendix E Convergence analysis of GeN
We provide the convergence analysis of GeN for multiple dimension, given that the 1-dimensional GeN is equivalent to the classic Newton-Raphson method. We work with the convex and Lipschitz conditions, the same conditions on which Prodigy and D-adaptation are proven to converge. We note that, similar to the proof of Newton’s method, we also need to bound the third order derivative and GeN can enjoy the quadratic convergence rate.
Theorem 1.
For convex and $G$-Lipschitz loss $L$, assuming $\mathbf{H}(x)\neq\bm{0}\forall x$ (i.e. no completely flat point), then GeN-GD in (3.2) with $\mathbf{g}\equiv\mathbf{G}$ gives
$\|\bm{w}_{t+1}-\bm{w}_{*}\|\leq\mathcal{M}\|\bm{w}_{t}-\bm{w}_{*}\|^{2}$
and
$L_{t}-L_{*}\leq G\|\bm{w}_{t}-\bm{w}_{*}\|\leq G\mathcal{M}^{2^{t}-1}\|\bm{w}_{0}-\bm{w}_{*}\|^{2^{t}}$
where $\bm{w}_{*}$ is the minimum of $L$, $\mathcal{M}=\sup_{x}||\kappa(x)||\cdot\sup_{x}1/||\mathbf{H}(x)||$, and $\kappa$ is the third order derivative of $L$. Furthermore, if $\mathcal{M}||\bm{w}_{0}-\bm{w}_{*}||<1$, then we have the quadratic convergence $\bm{w}_{t}\to\bm{w}_{*}$ as $t\to\infty$.
Proof.
Denote $\bm{e}_{t}=\bm{w}_{t}-\bm{w}_{*}$, then
$\bm{w}_{t+1}=\bm{w}_{t}-\frac{\mathbf{G}\mathbf{G}^{\top}\mathbf{G}}{\mathbf{G}^{\top}\mathbf{H}\mathbf{G}}\Big{|}_{\bm{w}_{t}}\Longrightarrow\bm{e}_{t+1}=\bm{e}_{t}-\frac{\mathbf{G}\mathbf{G}^{\top}\mathbf{G}}{\mathbf{G}^{\top}\mathbf{H}\mathbf{G}}\Big{|}_{\bm{e}_{t}+\bm{w}_{*}}$
Taylor expansion on the gradient at $\bm{w}_{t}$ leads to
$\mathbf{G}(\bm{w}_{*})=\mathbf{G}(\bm{w}_{t})-\mathbf{H}(\bm{w}_{t})\bm{e}_{t}+\kappa(\xi_{t})[\bm{e}_{t}]\bm{e}_{t}$
where $\mathbf{G}(\bm{w}_{*})=\bm{0}\in\mathbb{R}^{d}$ and $\kappa(\xi_{t})=\nabla^{3}L(\xi_{t})=\nabla\mathbf{H}(\xi_{t})\in\mathbb{R}^{d\times d\times d}$ is the third-order remainder. We have denoted the directional derivative $\kappa(x)[\bm{e}_{t}]=\lim_{h\to 0}\frac{\mathbf{H}(x+h\bm{e}_{t})-\mathbf{H}(x)}{h}\in\mathbb{R}^{d\times d}$.
Left-multiplying by $\mathbf{G}_{t}^{\top}=\mathbf{G}(\bm{w}_{t})^{\top}$ gives
$\mathbf{G}_{t}^{\top}\mathbf{H}_{t}\bm{e}_{t}-\mathbf{G}_{t}^{\top}\mathbf{G}_{t}=\mathbf{G}_{t}^{\top}\kappa(\xi_{t})[\bm{e}_{t}]\bm{e}_{t}$
Multiplying the generalized inverse $(\mathbf{G}_{t}^{\top}\mathbf{H}_{t})^{-1}$ gives
$\bm{e}_{t+1}=\bm{e}_{t}-\frac{\mathbf{G}_{t}\mathbf{G}_{t}^{\top}\mathbf{G}_{t}}{\mathbf{G}_{t}^{\top}\mathbf{H}_{t}\mathbf{G}_{t}}=\frac{\mathbf{G}_{t}\mathbf{G}_{t}^{\top}\kappa(\xi_{t})[\bm{e}_{t}]\bm{e}_{t}}{\mathbf{G}_{t}^{\top}\mathbf{H}_{t}\mathbf{G}_{t}}.$
By the matrix norm inequality, we obtain
$||\bm{e}_{t+1}||\leq\mathcal{M}||\bm{e}_{t}||^{2}$
i.e. the quadratic convergence in the parameter space. Furthermore, the Lipschitz condition allows the quadratic convergence in the loss space. ∎